A 6.25 kg bookbag is sitting on the ground 6.37 × 10 to the 6th power metersfeom the center pf the earth (m=5.972× 10 24th power kg)
What is the gravitational forse applie to the backpack?
F=ma divide your force from a) by the mass of backpack what is your aceleration? Why is this familiar
1
Expert's answer
2018-11-20T15:53:10-0500

a)

F=(6.25)(5.972∙ 〖10〗^24 )(6.67∙ 〖10〗^(-11) )/(6.37∙ 〖10〗^6 )^2 =61.4 N.

b)

a=(5.972∙ 〖10〗^24 )(6.67∙ 〖10〗^(-11) )/(6.37∙ 〖10〗^6 )^2 =9.8 m/s^2

a=g

It is well-known constant.